OpenStudy (anonymous):

your backyard: 3x^3-4x+8x^2-11 (length) 3x^2-6x^3+2 (width) neighbors backyard: 4-11x^2-17x (length) 7x^2-3x+9 (width) what is the difference between the perimeter of yours and your neighbors?

OpenStudy (anonymous):

they really stretch for these word problems, don't they?

OpenStudy (anonymous):

yes yes they do :-(((

OpenStudy (anonymous):

:-((( @saifoo.khan

OpenStudy (saifoo.khan):

Don't worry! Satellite will take care of everything!

OpenStudy (saifoo.khan):

Satellite is trolling today! D:

OpenStudy (anonymous):

can u help me with something to @saifoo.khan

OpenStudy (anonymous):

your perimeter \[2(3x^3-4x+8x^2-11) +2(3x^2-6x^3+2) \] \[=-6 x^3+22 x^2-8 x-18\]

OpenStudy (anonymous):

howd u do that??? @satellite73

OpenStudy (anonymous):

there perimeter \[2(4-11x^2-17x )+2(7x^2-3x+9)\] \[=-8 x^2-40 x+26\]

OpenStudy (anonymous):

thats only the perimeter for your yard @satellite73

OpenStudy (anonymous):

*their

OpenStudy (anonymous):

difference \[ -6 x^3+22 x^2-8 x-18-(-8 x^2-40 x+26)\] \[= -2 (3 x^3-15 x^2-16 x+22)\]